OpenStudy (anonymous):

given that the equation of a circle is x2 + y2-10x+4y +13=0 find its center and its radius

OpenStudy (anonymous):

12. Well, we have to complete the square. x^2 - 10x + 25 + y^2 + 4y + 4 + 13 = 25 + 4 (x - 5)^2 + (y + 2)^2 = 16 It's center, (5, -2), and it's radius is sqrt(16), or 4.